OpenStudy (anonymous):

Evaluate if a = -4, b = 5, and c = -3. -a-(-b)/c

OpenStudy (aravindg):

just apply BODMAS

OpenStudy (anonymous):

\[\frac{ -a-(-b) }{ c } \]

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

\[\frac{ -(-4)-(-5) }{ (-3) } \]

OpenStudy (anonymous):

-5.6

OpenStudy (anonymous):

....

OpenStudy (anonymous):

I did not get that.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

yeah show me ur work?

OpenStudy (anonymous):

BTW did you distribute the negative?

OpenStudy (anonymous):

Refer to the attached Mathematica solution.

OpenStudy (anonymous):

@some_someone \[4-\frac{5}{3}=\frac{7}{3} \]

OpenStudy (anonymous):

I got -3 as answer.

OpenStudy (anonymous):

@Directrix